Tech Staff Meeting Notes February 6, 2017
The top priority is always repair of down stations.
NM problem stations
ET problem stations
Duty Table updated June 2014
Fish Fry is tentatively scheduled for Friday May 5, Cinco de Mayo.
Hardware for ET network is entered. John and James will work together to load dataless.
Dave and James have started loading hardware for NM and AG.
Mitch will work with Kent to generate a station map for the public display.
GX620 at mttnworm needs to be visually checked. All others with "puffy caps" have been repaired. We have five spare motherboards.
BACM - spikes
BRNM - "
CPAR - "
TYAR - "
Strange artifacts on helis
RVAR see image
AG Misc Problems
X702 - keeps dropping out
Z41A is down
Low solar dropouts have several different sources though common problem is an extraordinarily overcast January. Other factors include mis-wiring, ISIS3 xmiters draw more power (so need more solar), AGM's have problems with low power charge, etc.
WYBT was vandalized for the second year in a row. Chris got it going temporarily. Need to revisit and harden. Done.
MORT is closed. Property owner plans to continue farming behind the levee and keep the road open. Will reinstall flood resistant MORT2 at a nicer spot.
There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.
I40 Bridge comms.
Jim Smith at NSMP is looking into radio options.
All nodes are due a firmware upgrade that will sufficiently change things such that the the initial upgrade and configure should be done at CERI, then swap hardware at the node. This will be done as machines are replaced.
DBB: Rebuild SFTN-MKTA portion of DBB.
Property at MKTA node will possibly be sold. Chris is looking into to it to make sure we can retain our presence.
Getting interference at the node causing spikes in CPAR, LVAR, LRAR, and TYAR.
AZTN/HDBR: Need to pull the 200' sensor at the weigh station. Probable water damage.
We have no dbb firewall spare. Steve will get a pair of replacements.
Flooding in the expansion space has become a problem. A Workorder was submitted and P-plant engineers were on-site to assess the problem. The most likely solution will be a retention wall and removal of material to make the foundation above grave. There are several open work orders. Jim will make a prioritized list and speak with zone supervisor.
Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob (MaggieValley), a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Compressor noise suppression is scheduled for the end of the month afterwhich a new noise study can be done.
Data taken from:
GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.
Obsidian and fanless PC for CVVA is being installed today. There's a possibility a true free field locations is available. John will investigate.
John's updated todo list .
Spikes at CVVA most likely originating at flat gap.
Improve RCGA AC power and perform maintenance work on shed.
Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are too tall to get a signal out. Looking for an alternate site.
HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established. Still waiting on completion of communcations to get on local internet. Its still on a cell modem; Randy Loop plans to install the internet modem next week. Hopefully after that we can retrieve the cell modem and get all channels to CERI in real-time.
Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.
GFM2 is good to go. It will be a reftek with S-13's. Do we need to take student volunteers to help with the GFM decommission?
WMTN pit needs to be moved due to erosion.
CPRT comm reconfig. John says, "CPRT is going through the Senior Center at Corryton (cort). We removed everything from svfd an everything but what is on the tower at gttn. cort is indeed using dynamic addressing but data is flowing. Greg Hill the director at the senior center is supposed to be contacting his IT person to see what can be done. I left a message for Greg last week but still haven't heard back. I want to get Greg's IT guy in direct contact with either Steve or James so they can figure it all out."
Site for newstation is permitted at a school in Armorel, AR just east of Blytheville. A $10k bid was submitted for installing the monument. It needs to be awarded before the end of the month.
CVMS has intermittent connection problems with wireless. Its hardwired for now but the wireless needs to be replaced to preserve electical isolation between the building and vault.
Z41A is online with AG. Vault leaks so did not install sm sensor. Need to fix the leak and install the sm vault inside the TA vault.
When the water leak at LCAR was repaired, was an accelerometer reinstalled?
Needs some clean up.
The FDL wiki is here: https://umwiki.memphis.edu/display/FDL/CERI+FDL
The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.
The seisnet wiki is at https://umwiki.memphis.edu/display/CERITech/CERI+Tech You will need your uuid and password.